OpenStudy (faman39):

Nadia took two practice exams to prepare for her college entrance exam. On the first practice exam, she scored 1,100 . After studying, her score increased by 30% . What was her score on the second practice exam?

OpenStudy (experimentx):

1100+30%of 1100 = 1100+0.3*1100 = 1100+330 = 1430

OpenStudy (wasiqss):

1430 is rite

Directrix (directrix):

1100 + 30% of 1100 = 1.3 (1100) = 1430